OpenStudy (linn99123):

NEED HELP FAST WILL GIVE MEDAL AND FAN Between what two consecutive integers does lie? A. 13 and 14 B. 12 and 13 C. 11 and 12 D. 10 and 11

OpenStudy (linn99123):



OpenStudy (anonymous):

i dont understand your question can you re word it

OpenStudy (noorik.):

between 12 and 13

OpenStudy (linn99123):

it says does the sqare root sigh and 50 sorry lol

OpenStudy (noorik.):

\[\sqrt{150} = 12.24\]

OpenStudy (anonymous):

care to explain how you got that @NooriK. ?

OpenStudy (noorik.):

because (12)^2=144 and (13)^2=169 and 150 between them and it's closer to 144 so i assumed root(150) would be 12.3 then used the calculator and i found out that it's 12.24

OpenStudy (linn99123):

umm thank u sir :D

OpenStudy (anonymous):

12 and 13 12.24 and my guess.... is he usd a calculatur

OpenStudy (anonymous):

good job mr. noori

OpenStudy (noorik.):

thanks